ROVIBRATIONAL CHARACTERIZATION OF X^{2}Σ^{+} ^{11}BH^{+} BY THE EXTRAPOLATION OF PHOTOSE-LECTED HIGH-RYDBERG SERIES IN ^{11}BH

C. RICARDO VITERI and ANDREW T. GILKISON^{a}, Department of Chemistry, Purdue University, West Lafayette, IN 47907; \underline{SCOTT J. RIXON} and EDWARD R. GRANT, Department of Chemistry, University of British Columbia, 6174 University Boulevard, Vancouver, BC, Canada V6T 1Z3.

Optical-optical-optical triple resonance spectroscopy of ^{11}BH isolates high-Rydberg states that form series converging to rotational state specific ionization potentials in the vibrational levels of ^{11}BH^{+} from v^{+} = 0 through 4. Limits defined by a comprehensive fit of these series to state-detailed thresholds yield rovibrational constants describing the X^{2}Σ^{+} state of 11BH^{+}. The data provide a first determination of the vibration-rotation interaction parameter α_{e} = 0.4821 cm^{−1} and a more accurate estimate of ω_{e} = 2526.58 cm^{−1} together with the higher-order anharmonic terms ω_{e}x_{e} = 61.98 cm^{−1} and ω_{e}y_{e} = −1.989 cm^{−1}. The deperturbation and global fit of series to state-detailed limits also yields a precise value of the adiabatic ionization potential of ^{11}BH of 79120.3(1) cm^{−1}, or 9.81033(1) eV. High precision is afforded here by the use of graphical analysis techniques, narrow bandwidth laser systems, and an analysis of newly observed, high-principal quantum number Rydberg states that conform well with a Hund’s case (d) electron-core coupling limit.

aPresent address: Schering-Plough, 2000 Galloping Hill Road, Kenilworth, NJ 07033-0530